• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name jpchen

    Log entry time 10:32:19 on July26,2006

    Entry number 175354

    keyword=day shift summary (first half)

    day shift summary:

    Shift workers: J. P. Chen (SL), G. Kumbartzki(TO), R. Sparks
    8:00 continue data taking for CSR test.
    Beam was not stable at the begining. First only 5 uA, then to 8 uA.

    Production run at HRS_L=120 degree, first kin (P0=100.97 MeV/c, with Pb-LH2
    8:35-8:50 changed to kin 2 P0=136.61 MeV/c, Pb-LH2
    9:05 move to slanted C
    9:35 moved to slanted Fe.
    9:45 MCC increasted our current to 20 uA since Hall C was down.
    Started BCM data logging for BCM calibration.
    10:03 stopped beam, stopped BCM data logging.
    CSR test completed.
    Requested hall to be open for restricted access for detector
    work (remove CSR calorimeter) and other work, which is expected
    last ~10 hours (Jack's estimation). Dave Meekins will do the target
    loop switch: switch LH2 from loop1 (which has a Pb foil inside) to loop2.

    run 1682, 8:05-8:35 kin_1, Pb/LH2, 410K
    run 1683, 8:50-,9:05 kin_2, Pb/LH2, 380K
    run 1684, 9:14-9:35 , kin_2, slanted C, 91K
    run 1685, 9:38-10:03 kin_2, slanted Fe, 230K, beam current increased
    during the run from 8 to 20 uA.